OpenStudy (anonymous):

Perform the following calculations to the correct number of significant figures. 41.4−2.59= 23.3+2.858+6.5+106.23= 18.4+54.60−5.904= 6.09−5.522=

OpenStudy (anonymous):

41.4-2.59=38.81 23.3+2.858+6.5+106.23=138.888 18.4+54.60-50904=67.096 6.09-5.522=0.568